# The Pólya-Tchebotarev problem with semiclassical external fields

### v1

@@ Appendix A Some technical lemmas on measures
Let $(\mu_{n})$ be a sequence of Borel probability measures on $\mathbb{C}$, and $\mu$ a Borel measure on $\mathbb{C}$, not necessarily a probability. The following statements are equivalent.
@@
A proof for this theorem can be found in [27, Theorem 3.2, Pg 52]. The first result we prove is a convergence statement for measures.
@@
Let $(\mu_{n})$ be a sequence of Borel probability measures on $\mathbb{C}$, and $\mu$ a Borel measure on $\mathbb{C}$, not necessarily a probability. Suppose that $\mu_{n}\to\mu$ vaguely. If $B$ is a Borel set for which $\mu(\partial B)=0$, then the sequence $({\mu_{n}}\raisebox{-1.29167pt}{$|$}_{B})$ converges vaguely to ${\mu}\raisebox{-1.29167pt}{$|$}_{B}$.
@@
Let $\mu$ be a finite Borel measure on $\mathbb{C}$. For each $R_{0}>0$ given, there exists $R>0$ for which
@@ Appendix B Quadratic differentials
On this section we briefly discuss relevant properties on quadratic differentials in $\overline{\mathbb{C}}$. For a full presentation on quadratic differentials, the interest reader is referred to the book by Strebel [34], and also Jensen’s chapter on Pommerenke’s book [24, Chapter 8].
A rational function $R$ defines on the Riemann sphere $\overline{\mathbb{C}}$ the quadratic differential $-R(z)\mathrm{d}z^{2}$. The critical points of $R$ are the zeroes and poles of $R$, including possibly the point $z=\infty$. A critical point $p\in\mathbb{C}$ has order
@@
A critical point $p$ is finite if $\operatorname{ord}(p)\geq-1$. An arc $\gamma$ is called an horizontal arc of $-R(z)\mathrm{d}z^{2}$ if, along $\gamma$,
@@
A maximal horizontal arc is called a horizontal trajectory, or simply a trajectory, of $-R(z)\mathrm{d}z^{2}$. A trajectory that ends at two finite critical points is called a critical trajectory. The local structure of trajectories is as follows:
@@
  • Each logarithmic singularity $w_{j}$ of $\Phi$ is a double pole of $-R(z)\mathrm{d}z^{2}$, and the residue of $\sqrt{-R(z)}$ at $w_{j}$ is exactly $\rho_{j}$. In particular, by (2.7), there are no closed trajectories around $p$;

### v2

@@ Appendix A Some technical lemmas on measures
Let $(\mu_{n})$ be a sequence of Borel probability measures on $\mathbb{C}$, and $\mu$ a Borel measure on $\mathbb{C}$, not necessarily a probability measure. The following statements are equivalent.
@@
A proof for this theorem can be found in [32, Theorem 3.2, pg 52]. The first result we prove is a convergence statement for measures.
@@
Let $(\mu_{n})$ be a sequence of Borel probability measures on $\mathbb{C}$, and $\mu$ a Borel measure on $\mathbb{C}$, not necessarily a probability measure. Suppose that $\mu_{n}\to\mu$ vaguely. If $B$ is a Borel set for which $\mu(\partial B)=0$, then the sequence $({\mu_{n}}\raisebox{-1.29167pt}{$|$}_{B})$ converges vaguely to ${\mu}\raisebox{-1.29167pt}{$|$}_{B}$.
@@
Let $\mu$ be a finite Borel measure on $\mathbb{C}$. For each $R_{0}>0$ given, there exists $R>R_{0}$ for which
@@ Appendix B Quadratic differentials
On this section we briefly discuss relevant properties on quadratic differentials in $\overline{\mathbb{C}}$. For a full presentation on quadratic differentials, the interest reader is referred to the book by Strebel [39], and also Jensen’s chapter on Pommerenke’s book [29, Chapter 8].
A rational function $R$ defines on the Riemann sphere $\overline{\mathbb{C}}$ the quadratic differential $-R(z)\mathrm{d}z^{2}$.
The critical points of $R$ are the zeroes and poles of $R$, including possibly the point $z=\infty$. A critical point $p\in\mathbb{C}$ has order
@@
A critical point $p$ is finite if $\operatorname{ord}(p)\geq-1$.
An arc $\gamma$ is called an horizontal arc of $-R(z)\mathrm{d}z^{2}$ if, along $\gamma$,
@@
A maximal horizontal arc is called a horizontal trajectory, or simply a trajectory, of $-R(z)\mathrm{d}z^{2}$. A trajectory that ends at two finite critical points is called a critical trajectory.
The local structure of trajectories is as follows:
@@
  • Each $w_{j}$ that is different from each $z_{i}$ is a double pole of $-R(z)\mathrm{d}z^{2}$, and the residue of $\sqrt{-R(z)}$ at $w_{j}$ is exactly $\rho_{j}$.